# Conceptualization, Data Curation, Formal Analysis, Methodology, Software, Writing – original draft

### v1

@@ A.1 General Flight Information Prompt
From the general flight information, which includes multiple data types as described in Table 2, we convert the data into a format compatible with language models using 10 distinct prompt templates. Examples of these prompt formats are illustrated in Figure 9. These templates convert structured data into natural language, aligning better with LLM pretraining knowledge. Moreover, using multiple formats improves prompt generalization of the trainable components by training the model with varied expressions of the same content.
@@
General Flight Information Prompt Format Example On {date}, a {haul}-haul flight by {airline_name_english} was scheduled to arrive at {dest_name_english} ({dest_code_icao}/{dest_code_iata}) at {sched_time_utc} UTC. Flight ID: {callsign_code_icao}/{callsign_code_iata}. The aircraft departed from {dep_name_english} ({dep_code_icao}, lat: {dep_lat}, lon: {dep_lon}, alt: {dep_altitude} ft) and was destined for {dest_name_english}, located at lat: {dest_lat}, lon: {dest_lon}, alt: {dest_altitude} ft. The aircraft was a {aircraft_type} with registration {aircraft_registration}, and it belonged to the {wake_turbulence_cat} wake turbulence category. The aircraft was expected to enter monitored airspace around {actual_entry_time}. Day of week: {day_of_week}, distance: {distance} km.
Figure 9: Examples of Prompt Format for General Flight Information Description
@@ A.2 Weather Prompt
The weather prompt $P^{\text{wx}}_{t}$ is constructed from the METAR and TAF data queried at time $t$ retrieved during dataset preparation. These coded weather reports are parsed and reformatted into natural language prompts following a predefined structure. An example of the formatted weather prompt is shown in Figure 10.
Weather Information Prompt Format Example METAR in effect: 202202272230 METAR RKSI 272230Z 14006KT 5000 BR NSC 01/M00 Q1023 NOSIG= TAF in effect: 202202271700 TAF RKSI 271700Z 2718/2824 21006KT CAVOK TN01/2722Z TX08/2806Z TN03/2820Z BECMG 2721/2722 14006KT 4000 BR NSC BECMG 2800/2801 21006KT 6000 NSW BECMG 2814/2815 14010KT BECMG 2817/2818 17006KT 3500 -RA FEW010 BKN025 OVC100 TEMPO 2822/2824 29006KT 0800 DZ FG BKN005
@@ A.3 NOTAM Prompt
The NOTAM prompt $P^{no}_{t}$ is constructed from all active NOTAMs at time $t$, capturing time-specific operational constraints. Instead of decoding aeronautical codes, we maintain the raw text and format it using a fixed template (Figure 11), similar to the weather information prompt.
Special Notices Prompt Format Example Active NOTAMs: REF AIP SUP 31/22, A-CDM PHASE 2 TRIAL OPERATION PROGRESSIVE STEPS 2 AND 3 PERIOD CHG AS FLW 1. STEP 2 : FROM 1600 UTC 29 JUN 2022 TO 1600 UTC 11 JAN 2023 2. STEP 3 : FROM 1600 UTC 11 JAN 2023 TO 1600 UTC 27 JUN 2023) TEMP OBST(CRANES) ERECTED AS FLW 1. PSN : 372806.54N 1262700.83E, 372740.28N 1262724.18E 2. RADIUS : 10M 3. HGT : 20.24M(66.4FT) AMSL) TEMP OBST(CRANES) ERECTED AS FLW 1. PSN : 372812.81N 1262551.43E, 372815.70N 1262553.51E, 372814.91N 1262557.46E, 372812.15N 1262552.41E 2. RADIUS : 40M 3. HGT : 60.39M(198.13FT) AMSL) TDWR(TERMINAL DOPPLER WEATHER RADAR) U/S DUE TO WIP)
@@ A.4 Static Guiding Prompt for Trajectories
Three types of trajectories, $X^{f}_{i,t}$, $X^{a}_{i,t}$, and $X^{p}_{i,t}$, represent distinct aspects of the airspace situation surrounding the flight of interest. To help the model distinguish and reason over these inputs, we introduce a structured prompt that explicitly labels each trajectory group. The guiding prompt segments specify which input embeddings correspond to each trajectory type, thereby helping the model distinguish their distinct semantic roles.
@@
The static prompts (Figure 12) will be tokenized and embedded in parallel with the trajectory embeddings obtained through a trajectory representation model and a cross-modality adaptation network. The instance-level representations of the trajectories are inserted between the static prompt segments, resulting in a unified language-time series embedding that describes the airspace. This design mirrors how human operators visualize dynamic traffic, thereby enabling effective modeling of airspace conditions without dependence on tasks such as traffic count, traffic classification, capacity assessment, or conflict detection.
@@ Appendix B Full Model Comparison Results
Table 5: Model Performance Metrics for Q1-Q2 (January–June)
@@
Table 6: Model Performance Metrics for Q3-Q4 (July–December)
@@ Appendix C Full Context Analysis Results
Table 7: Model Performance for Q1 (January–March)
@@
Table 8: Model Performance for Q2 (April–June)
@@
Table 9: Model Performance for Q3 (July–September)
@@
Table 10: Model Performance for Q4 (October-December)

### v2

@@ A.1 General Flight Information Prompt
The general flight information (Table 2) is converted into an LLM-compatible format using ten prompt templates, with an example illustrated in Figure 9. These templates transform structured data into natural language consistent with LLM pretraining, while multiple formats enhance generalization by exposing the model to varied expressions of the same content.
@@
Figure 9: Example of Prompt Format for General Flight Information Description
@@ A.2 Weather Prompt
The weather prompt $P^{\text{wx}}_{t}$ is constructed from the METAR and TAF data queried at time $t$, parsed and reformatted into a predefined structure. An example is shown in Figure 10.
Weather Information Prompt Example METAR in effect: 202202272230 METAR RKSI 272230Z 14006KT 5000 BR NSC 01/M00 Q1023 NOSIG= TAF in effect: 202202271700 TAF RKSI 271700Z 2718/2824 21006KT CAVOK TN01/2722Z TX08/2806Z TN03/2820Z BECMG 2721/2722 14006KT 4000 BR NSC BECMG 2800/2801 21006KT 6000 NSW BECMG 2814/2815 14010KT BECMG 2817/2818 17006KT 3500 -RA FEW010 BKN025 OVC100 TEMPO 2822/2824 29006KT 0800 DZ FG BKN005
@@ A.3 NOTAM Prompt
The NOTAM prompt $P^{no}_{t}$ is constructed from all active NOTAMs at time $t$. We maintain the raw text and format it using a fixed template (Figure 11).
Special Notices Prompt Example Active NOTAMs: REF AIP SUP 31/22, A-CDM PHASE 2 TRIAL OPERATION PROGRESSIVE STEPS 2 AND 3 PERIOD CHG AS FLW 1. STEP 2 : FROM 1600 UTC 29 JUN 2022 TO 1600 UTC 11 JAN 2023 2. STEP 3 : FROM 1600 UTC 11 JAN 2023 TO 1600 UTC 27 JUN 2023) TEMP OBST(CRANES) ERECTED AS FLW 1. PSN : 372806.54N 1262700.83E, 372740.28N 1262724.18E 2. RADIUS : 10M 3. HGT : 20.24M(66.4FT) AMSL) TEMP OBST(CRANES) ERECTED AS FLW 1. PSN : 372812.81N 1262551.43E, 372815.70N 1262553.51E, 372814.91N 1262557.46E, 372812.15N 1262552.41E 2. RADIUS : 40M 3. HGT : 60.39M(198.13FT) AMSL) TDWR(TERMINAL DOPPLER WEATHER RADAR) U/S DUE TO WIP)
@@ A.4 Static Guiding Prompt for Trajectories
Three types of trajectories, $X^{f}_{i,t}$, $X^{a}_{i,t}$, and $X^{p}_{i,t}$, capture different aspects of the airspace condition. To help the model reason over them, structured guiding prompts label each trajectory group, clarifying their distinct semantic roles.
@@
The static prompts (Figure 12) are tokenized and embedded alongside trajectory embeddings from the representation and adaptation networks. Instance-level trajectory representations are inserted between prompt segments, forming a trajectory-informed sequence (Equation 22).
@@ Appendix B Full Model Comparison Results
Table 5: Comparison Study of LLM backbone Performance for Q1-Q2 (January-June)
@@
Table 6: Comparison Study of LLM backbone Performance for Q3-Q4 (July-December)
@@ Appendix C Full Context Analysis Results
Table 7: Performance of Meta-LLaMA-3.2-1B-Instruct under Different Context Removal Settings for Q1 (January-March)
@@
Table 8: Performance of Meta-LLaMA-3.2-1B-Instruct under Different Context Removal Settings for Q2 (April-June)
@@
Table 9: Performance of Meta-LLaMA-3.2-1B-Instruct under Different Context Removal Settings for Q3 (July-September)
@@
Table 10: Performance of Meta-LLaMA-3.2-1B-Instruct under Different Context Removal Settings for Q4 (October-December)
@@
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT and Grammarly to improve the language and readability. After using these tools, the authors reviewed and edited the content as needed and take full responsibility for the final version of the publication.